OpenStudy (anonymous):

find the inverse of the function and graph the function and its inverse on the same set of axes

OpenStudy (anonymous):

\[f(x)=3^{x+2}-4\]

OpenStudy (anonymous):

@hartnn ??

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

i know i have to take the log of it to get the exponent right ?

ganeshie8 (ganeshie8):

yup! keep going...

ganeshie8 (ganeshie8):

solve for x first

ganeshie8 (ganeshie8):

\(f(x)=3^{x+2}-4\) \(x = ?\)

OpenStudy (anonymous):

??

OpenStudy (anonymous):

add 4 ?

ganeshie8 (ganeshie8):

yes add 4 both sides

ganeshie8 (ganeshie8):

\(f(x) = 3^{x+2}-4\) add 4 both sides \(f(x)+4 = 3^{x+2}\) take \(log_3\) both sides \(\log_3 (f(x)+4) = \log_3(3^{x+2})\) \(\log_3 (f(x)+4) = (x+2) \log_3(3)\) \(\log_3 (f(x)+4) = x+2 \) subtract 2 both sides \(\log_3 (f(x)+4) -2 = x \) switch x and f(x) \(\log_3 (x+4) -2 = y\)

ganeshie8 (ganeshie8):

thats ur inverse funciton

ganeshie8 (ganeshie8):

sketch both the functions on one graph paper

OpenStudy (ikram002p):



OpenStudy (ikram002p):